### v1

@@ 5 Comparison of standard and multiprecision Schwarz methods
(31) $\tilde{A}_{i}^{-1}-A_{i}^{-1}=\mathcal{E}_{i}A_{i}^{-1}.$
As a result, if we assume******This assumption is analogous to (16). In fact, the derivations requiring (16) can be carried out analogously even if we assume (32) instead of (16) but the derivation becomes somewhat more lengthy.
(32) $\|\mathcal{A}_{i}^{-1}\mathtt{F}_{i}\|\leq\epsilon<\frac{1}{2},$
@@
(33) $\|\tilde{A}_{i}^{-1}-A_{i}^{-1}\|=2\epsilon\kappa\left(D_{i}^{(c)}\right)\|A_{i}^{-1}\|,$
@@
(34) $\tilde{M}_{\star}^{-1}A=M_{\star}^{-1}A+E_{\star}\quad\mathrm{where}\quad E_{\star}:=\begin{cases}\theta\sum\limits_{i=1}^{p}R_{i}^{T}\mathcal{E}_{i}A_{i}^{-1}R_{i}A,&\mathrm{for\;(damped)\;AS},\\ \sum\limits_{i=1}^{p}\bar{R}_{i}^{T}\mathcal{E}_{i}A_{i}^{-1}R_{i}A,&\mathrm{for\;RAS}.\end{cases}$
@@
(35) $E_{\star}=\begin{cases}AS,\theta\,:\quad&\theta\sum\limits_{i=1}^{p}\Pi_{i}^{T}\begin{bmatrix}\mathcal{E}_{i}&0\\ 0&0\end{bmatrix}\begin{bmatrix}I_{N_{i}}&A_{i}^{-1}K_{i}\\ 0&0\end{bmatrix}\Pi_{i},\\ RAS\,:\quad&\sum\limits_{i=1}^{p}\overline{\Pi}_{i}^{T}\begin{bmatrix}\begin{bmatrix}(\mathcal{E}_{i})_{1:\bar{N}_{i},:}\\ 0_{\bar{N}_{i}+1:N_{i},:}\end{bmatrix}&0\\ 0&0\end{bmatrix}\begin{bmatrix}I_{N_{i}}&A_{i}^{-1}K_{i}\\ 0&0\end{bmatrix}\Pi_{i},\end{cases}$
@@
(36) $E_{AS,\theta}=\theta\Pi^{T}\begin{bmatrix}\mathcal{E}_{1}&0\\ 0&\mathcal{E}_{2}\end{bmatrix}\begin{bmatrix}I_{N_{1}}&A_{1}^{-1}K_{1}\\ A_{2}^{-1}K_{2}&I_{N_{2}}\end{bmatrix}Z\Pi,$
@@
We notice that this way we managed to factor the error matrix $E_{AS,\theta}^{(k)}$ so that the first term carries the multiprecision error while the second term carries the Schwarz-method structure. As a result, bounding the norm of $E_{AS,\theta}^{(k)}$ and $E_{AS,\theta}$ becomes easier*†*†*†In many areas of interest it is often more suitable to bound the norm of the iteration matrix (and hence of the error) over two or more iterations due to the nature of the underlying PDE analysis, see, e.g., [23, 26]. Further research in this direction might be useful here as well.. Assuming (32) for both $i=1,2$ and noticing that $\|Z\|\leq 2$ we observe that
(37) $\|E_{AS,\theta}\|\leq 2\epsilon\kappa\left(D_{i}^{(c)}\right)\|M^{-1}_{AS,\theta}A\|.$
@@
A natural next step would be to carry out this reformulation also for the iteration matrix $\tilde{T}_{AS,\theta}$ as it is its spectral radius that asymptotically governs the convergence. However, since the spectral radius is not sub-additive or even stable with respect to perturbations, this wouldn’t be directly useful for quantifying the slow-down of the Schwarz method convergence, unless we consider a more specific situation. An extreme example is the case when the rounding is in nature only scalar (highlighting the “nicest” case included in the above setting). That is, there exist a scalar $\alpha$ such that $\mathcal{A}_{i}=\alpha\mathtt{A}_{i}$ where $\tilde{\mathtt{A}}_{i}=\mathtt{A}_{i}$, i.e., the subdomain matrices are scalar multiples of a matrix that can be stored “exactly”*‡*‡*‡Here “exactly” means to the same precision we store the solution. Also, notice that verifying (12) becomes trivial. Notice that such problems arise, e.g., when discretizing “nice” Poisson-like problems with finite differences so that $\alpha=1/h^{2}$. in the considered precision $u_{\ell}$. Then
(38) $\mathtt{F}_{i}=\left(\tilde{\alpha}-\alpha\right)\mathcal{A}_{i},$
@@
A direct calculation then gives the formulasMany of the following calculations are similar to the ones presented in [26, Section 3.2] for the modified restricted Schwarz method.
@@
(40) $\frac{\|\mathbf{r}_{k}\|}{\|\mathbf{r}_{0}\|}\leq\frac{L_{\delta}}{2\pi\delta}\;\min\limits_{\begin{subarray}{c}\deg(\varphi)\leq k\\ \varphi(0)=1\end{subarray}}\;\max\limits_{z\in\sigma_{\delta}(\mathrm{X})}|\varphi(z)|,$
@@
(41) $\frac{\|\bm{\rho}_{k}\|}{\|\mathbf{b}\|}\leq\left(1+\frac{\epsilon}{\delta-\epsilon}\right)\frac{L_{\delta}}{2\pi\delta}\;\min\limits_{\begin{subarray}{c}\deg(\varphi)\leq k\\ \varphi(0)=1\end{subarray}}\;\max\limits_{z\in\sigma_{\delta}(\mathrm{M^{-1}_{\star}A})}|\varphi(z)|.$

### v2

@@ 5 Comparison of standard and multiprecision Schwarz methods
$\tilde{A}_{i}^{-1}-A_{i}^{-1}=\mathcal{E}_{i}A_{i}^{-1}.$ (31)
As a result, if we assume [^12] This assumption is analogous to (16). In fact, the derivations requiring (16) can be carried out analogously even if we assume (32) instead of (16) but the derivation becomes somewhat more lengthy.
$\|\mathcal{A}_{i}^{-1}\mathtt{F}_{i}\|\leq\epsilon<\frac{1}{2},$ (32)
@@
$\|\tilde{A}_{i}^{-1}-A_{i}^{-1}\|=2\epsilon\kappa\left(D_{i}^{(c)}\right)\|A_{i}^{-1}\|,$ (33)
@@
$\tilde{M}_{\star}^{-1}A=M_{\star}^{-1}A+E_{\star}\quad\mathrm{where}\quad E_{\star}:=\begin{cases}\theta\sum\limits_{i=1}^{p}R_{i}^{T}\mathcal{E}_{i}A_{i}^{-1}R_{i}A,&\mathrm{for\;(damped)\;AS},\\ \sum\limits_{i=1}^{p}\bar{R}_{i}^{T}\mathcal{E}_{i}A_{i}^{-1}R_{i}A,&\mathrm{for\;RAS}.\end{cases}$ (34)
@@
$E_{\star}=\begin{cases}AS,\theta\,:\quad&\theta\sum\limits_{i=1}^{p}\Pi_{i}^{T}\begin{bmatrix}\mathcal{E}_{i}&0\\ 0&0\end{bmatrix}\begin{bmatrix}I_{N_{i}}&A_{i}^{-1}K_{i}\\ 0&0\end{bmatrix}\Pi_{i},\\ RAS\,:\quad&\sum\limits_{i=1}^{p}\overline{\Pi}_{i}^{T}\begin{bmatrix}\begin{bmatrix}(\mathcal{E}_{i})_{1:\bar{N}_{i},:}\\ 0_{\bar{N}_{i}+1:N_{i},:}\end{bmatrix}&0\\ 0&0\end{bmatrix}\begin{bmatrix}I_{N_{i}}&A_{i}^{-1}K_{i}\\ 0&0\end{bmatrix}\Pi_{i},\end{cases}$ (35)
@@
$E_{AS,\theta}=\theta\Pi^{T}\begin{bmatrix}\mathcal{E}_{1}&0\\ 0&\mathcal{E}_{2}\end{bmatrix}\begin{bmatrix}I_{N_{1}}&A_{1}^{-1}K_{1}\\ A_{2}^{-1}K_{2}&I_{N_{2}}\end{bmatrix}Z\Pi,$ (36)
@@
We notice that this way we managed to factor the error matrix $E_{AS,\theta}^{(k)}$ so that the first term carries the multiprecision error while the second term carries the Schwarz-method structure. As a result, bounding the norm of $E_{AS,\theta}^{(k)}$ and $E_{AS,\theta}$ becomes easier [^13] In many areas of interest it is often more suitable to bound the norm of the iteration matrix (and hence of the error) over two or more iterations due to the nature of the underlying PDE analysis, see, e.g., [23, 26]. Further research in this direction might be useful here as well.. Assuming (32) for both $i=1,2$ and noticing that $\|Z\|\leq 2$ we observe that
$\|E_{AS,\theta}\|\leq 2\epsilon\kappa\left(D_{i}^{(c)}\right)\|M^{-1}_{AS,\theta}A\|.$ (37)
@@
A natural next step would be to carry out this reformulation also for the iteration matrix $\tilde{T}_{AS,\theta}$ as it is its spectral radius that asymptotically governs the convergence. However, since the spectral radius is not sub-additive or even stable with respect to perturbations, this wouldn’t be directly useful for quantifying the slow-down of the Schwarz method convergence, unless we consider a more specific situation. An extreme example is the case when the rounding is in nature only scalar (highlighting the “nicest” case included in the above setting). That is, there exist a scalar $\alpha$ such that $\mathcal{A}_{i}=\alpha\mathtt{A}_{i}$ where $\tilde{\mathtt{A}}_{i}=\mathtt{A}_{i}$, i.e., the subdomain matrices are scalar multiples of a matrix that can be stored “exactly” [^14] Here “exactly” means to the same precision we store the solution. Also, notice that verifying (12) becomes trivial. Notice that such problems arise, e.g., when discretizing “nice” Poisson-like problems with finite differences so that $\alpha=1/h^{2}$. in the considered precision $u_{\ell}$. Then
$\mathtt{F}_{i}=\left(\tilde{\alpha}-\alpha\right)\mathcal{A}_{i},$ (38)
@@
A direct calculation then gives the formulas [^15] Many of the following calculations are similar to the ones presented in [26, Section 3.2] for the modified restricted Schwarz method.
@@
$\frac{\|\mathbf{r}_{k}\|}{\|\mathbf{r}_{0}\|}\leq\frac{L_{\delta}}{2\pi\delta}\;\min\limits_{\begin{subarray}{c}\deg(\varphi)\leq k\\ \varphi(0)=1\end{subarray}}\;\max\limits_{z\in\sigma_{\delta}(\mathrm{X})}|\varphi(z)|,$ (40)
@@
$\frac{\|\bm{\rho}_{k}\|}{\|\mathbf{b}\|}\leq\left(1+\frac{\epsilon}{\delta-\epsilon}\right)\frac{L_{\delta}}{2\pi\delta}\;\min\limits_{\begin{subarray}{c}\deg(\varphi)\leq k\\ \varphi(0)=1\end{subarray}}\;\max\limits_{z\in\sigma_{\delta}(\mathrm{M^{-1}_{\star}A})}|\varphi(z)|.$ (41)